OpenStudy (anonymous):

Plastic parts produced by an injection-molding operation are checked for conformances to specifications. Each tool contains 12 cavities in which parts are produced, and these parts fall into a conveyor when the press opens. An inspector chooses three parts from the 12 at random. Two cavities are affected by a temperature malfunction that results in parts that do not conform to specifications. a) What is the probability that the inspector finds exactly one nonconforming part? b) What is the probability that the inspector finds at least one nonconforming part?

OpenStudy (anonymous):

a. There are 12c3 = 220 ways for the inspector to pick 3 parts to check. When exactly one of these three is bad: there are 2 ways to pick the bad part and 10c2 = 45 ways to pick good parts. 2 * 45/220 = 9/22. b. The probability that the inspector checks only good parts is 10c3/12c3 = 120/220 = 6/11. Therefore, the probability that he finds at least one bad is 1 - 6/11 = 5/11.

OpenStudy (anonymous):

Thanks, appreciated

OpenStudy (anonymous):

You're welcome!! Xo!